### v2

# entanglement.bib

Abstract.
Spectral properties of bounded linear operators play a crucial role in several areas of mathematics and physics. For each self-adjoint, trace-class operator $O$ we define a set $\Lambda_{n}\subset\mathbb{R}$, and we show that it converges to the spectrum of $O$ in the Hausdorff metric under mild conditions. Our set $\Lambda_{n}$ only depends on the first $n$ moments of $O$. We show that it can be effectively calculated for physically relevant operators, and it approximates the spectrum well.
We prove that using the above method we can converge to the minimal and maximal eigenvalues with super-exponential speed.
We also construct monotone increasing lower bounds $q_{n}$ for the minimal eigenvalue (or decreasing upper bounds for the maximal eigenvalue). This sequence only depends on the moments of $O$ and a concrete upper estimate of its $1$-norm; we also demonstrate that $q_{n}$ can be effectively calculated for a large class of physically relevant operators. This rigorous lower bound $q_{n}$ tends to the minimal eigenvalue with super-exponential speed provided that $O$ is not positive semidefinite. As a by-product, we obtain computable upper bounds for the $1$-norm of $O$, too.
Key words and phrases:
spectral theory, self-adjoint, trace-class, integral operators, density operators, open quantum systems, quantum theory, positive semidefinite, Schatten–von Neumann norm, Newton’s identities, Faddeev–LeVerrier algorithm, entanglement
2020 Mathematics Subject Classification:
46N50, 47A75, 47B65, 47G10, 81Q10
1. Introduction
1.1. History and motivation
Self-adjoint trace-class linear operators on infinite dimensional separable Hilbert spaces played an inevitable role in the foundation of quantum mechanics [Neumann, Weyl, Wigner, Husimi, Moyal]. In particular, density operators [Neumann] have found many applications in quantum chemistry, statistical mechanics, quantum optics, quantum information theory, and the theory of open quantum systems [Hillery, Lee, Schleich, Weinbub, Salazar, book1, HPZ, H-Y]. These operators occur in mathematical physics [grothendieck1966produits, dubinsky1979structure, wong1979schwartz], and they recently appeared in the modern physical description of our universe [Fahn_2023, hsiang2024graviton].
Finding the spectrum of linear operators, or deciding whether they are at least positive semidefinite is an important and delicate problem: positivity preservation during the time evolution is a crucial question in the dynamics of open quantum systems [Gnutzmann, Gosson, Newton, BLH, HBCSCS, HCSB]. Quantum entanglement problems still pose hard challenges to mathematicians and theoretical physicists [EPR, Bell, HORODECKI19961, Open]; they are also intimately connected to positivity via the Peres–Horodecki criterion [HORODECKI19961, Perescrit]. Describing entanglement in a quantitative way is a very important task in the theory of quantum computing, and quantum communication. Therefore, it is useful to know the magnitude of the negativity111the sum of the absolute values of the negative eigenvalues of a density operator after the partial transpose [Duan, Ruggiero, Kumari, Kao_2016]. While the spectrum can be calculated relatively easily in finite dimensional Hilbert spaces [separable, Eisert, Vidal, Plenio], for infinite dimensional non-Gaussian operators it is still a hard problem [Lami_2023, Pirandola, Vogel, Miki]. Infinite dimensional trace-class integral operators arise in the modeling of quantum optical channels, the theory of distributed quantum communication, and the quantum internet [Sharma2023, pathak2013elements, Kao_2016, Rohde_2021].
The spectrum of a trace-class operator determines a lot of important quantities, arguably the most important and difficult ones are the entropies, for example the Rényi or von Neumann entropy, see [Shirokov2010]. Entropy is a crucial notion in the areas of quantum information theory [nielsen_chuang_2010], quantum communication [Sharma2023], quantum computation [pathak2013elements], quantum cryptography [Yao_22] just to name a few. The eigenvalues and kernels of self-adjoint trace class integral operators has also many applications in the area of machine learning, see e. g. [DBL] and references therein. As the eigenvalues and eigenvectors of a density operator encode all information about the physical subsystem it describes [Neumann, Neumann1927, Lloyd2014], approximating the spectrum has far-reaching physical applications and mathematical relevance.
One of the main challenges in quantum technology is to filter out quantum noise stemming from the imprecise quantum physical measurements and other external circumstances [clerk2010introduction]. Mathematical and physical properties of quantum noise and quantum decoherence are only understood in simple models so far [chang2015quantum, beny2009quantum]. Understanding the dynamics of entanglement between two coupled Brownian oscillators (considered as an open quantum system) is a notoriously hard problem [entang_Hu]. Here time evolution is described by master equations which are very hard to solve analytically, and the problem is usually formalized in infinite dimensional Hilbert spaces, which also adds to the challenge. The Gaussian case can be solved exactly under certain conditions [entang_Hu], but the non-Gaussian case is still mostly unexplored. Here approximating the eigenvalues yields estimate for the fidelity [nielsen_chuang_2010] and for various quantum entropies [ohya2004quantum].
The maximal eigenvalue also has a meaning in statistical physics as follows. We consider a Hamiltonian operator $\hat{H}$ of a physical system, whose energy spectrum is supposed to have a lower bound. The eigenstate with the lowest energy is called the ground state. In many cases, after a physical measurement, the examined physical system can be found in the ground state with larger probability than in any other state. Also, the largest probability equals the maximal eigenvalue of the statistical physical density operator. For example, in case of canonical ensembles in statistical quantum physics, the equilibrium state is characterized by a density operator $\hat{\rho}=e^{-\beta\hat{H}}/\operatorname{Tr}\big{\{}e^{-\beta\hat{H}}\big{\}}$, see [reif2009fundamentals]. Here the maximal eigenvalue equals the probability of being in the ground state; this probability is an important quantity e. g. in the quantum thermodynamic description of physical systems [diosi2011short].
The minimal eigenvalue of a self-adjoint operator, if negative, is a quantitative measure how far the operator is from being positive semidefinite.
Estimating the Schatten–von Neumann $1$-norm of trace-class integral operators plays an important role in the mathematics of anharmonic oscillators [DELGADO20211] and areas of harmonic analysis related to physics [Catana].
Our main examples will be polynomial Gaussian operators, see Definition 2.1. For the physical relevance of polynomial Gaussian operators and for qualitative results about their positivity the readers might consult [balka2024positivity].
1.2. Summary of our results and the structure of the paper
Let $\mathcal{H}$ be a separable, complex Hilbert space and let $O\colon\mathcal{H}\to\mathcal{H}$ be a self-adjoint, trace-class linear operator. Then $O$ has only countably many real eigenvalues, let us enumerate them with multiplicities as $\{\lambda_{i}\}_{i\geq 0}$. As $O$ is trace-class, we have $\sum_{i\geq 0}|\lambda_{i}|<\infty$. If there are only finitely many eigenvalues $\lambda_{0},\dots,\lambda_{m}$ then define $\lambda_{i}=0$ for each integer $i>m$.
Following [Newton] we define symmetric functions of the eigenvalues $e_{k}$ such that $e_{0}=1$ and for $k\geq 1$ we have
$e_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{0\leq i_{1}<\dots<i_{k}}\lambda_{i_{1}}\cdots\lambda_{i_{k}};$
we can calculate $e_{k}$ from the moments of $O$, see Subsection 3.1. In Section 3 we introduce the finite set $\Lambda_{n}\subset\mathbb{R}$ as
$\Lambda_{n}\stackrel{{\scriptstyle\text{def}}}{{=}}\left\{-\frac{1}{x}:x\in\mathbb{R}\text{ and }\sum_{k=0}^{n}e_{k}x^{k}=0\right\}.$
In order to find $\Lambda_{n}$, we only need to calculate the above mentioned quantities $e_{k}$ for $k=0,\dots,n$. We prove in Theorem 3.4 that $\Lambda_{n}$ converges to the closure of the set of non-zero eigenvalues in the Hausdorff metric, see Subsection 3.4. In Subsection 3.2 we demonstrate by a Gaussian example that $\Lambda_{n}$ is very close to the set of eigenvalues even for $n=10$; in this case the exact spectrum is known and can be compared to $\Lambda_{n}$. In Subsection 3.3 we calculate $\Lambda_{10}$ for a non-Gaussian example as well, where the exact spectrum is unknown.
In the latter sections we approximate the extremal eigenvalues and verify faster convergence. We only consider the case of the minimal eigenvalue, since replacing the operator $O$ with $-O$ handles the maximal eigenvalue as well. Let us define
$\lambda_{\min}\stackrel{{\scriptstyle\text{def}}}{{=}}\min\{0,\lambda_{i}:i\geq 0\},$
[…]
In Section 4 we approximate $\lambda_{\min}$ by the sequence $\{q_{n,0}\}_{n\geq 1}$, where
$q_{n,0}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{-1}{\min\left\{x>0:\sum_{k=0}^{n}e_{k}x^{k}=0\right\}}.$
In Theorem 4.3 we prove that $q_{n,0}\to\lambda_{\min}$ with super-exponential speed:
$|q_{n,0}-\lambda_{\min}|\leq\exp\left[-n\log(n)+\mathcal{O}(n)\right]$
under the slight technical condition that $\lambda_{\min}$ is negative and has multiplicity one.
In Section 6 we assume that $c>0$ is a number satisfying
(1.1) $\sum_{i\geq 0}|\lambda_{i}|\leq c.$
Then we can construct and calculate a sequence $\{q_{n,c}\}_{n\geq 0}$ such that
  1. (i)
    $q_{n,c}$ is monotone increasing and $q_{n,c}\to\lambda_{\min}$ as $n\to\infty$;
  2. (ii)
    $q_{n,c}\approx-\frac{c}{n}$ if $O$ is positive semidefinite;
  3. (iii)
    $q_{n,c}\to\lambda_{\min}$ with super-exponential speed if $O$ is not positive semidefinite.
In Section 5 we demonstrate that we can effectively find values $c$ according to (1.1) for a large class of integral operators having polynomial Gaussian kernels, see Definition 2.1 for the precise notion. Finding an upper bound for the Schatten–von Neumann $1$-norm $\sum_{i\geq 0}|\lambda_{i}|$ is a hard problem in itself. We do this by writing $O$ as the product of two, typically not self-adjoint Hilbert–Schmidt integral operators in different ways, and we apply Hölder’s inequality (5.3); finally, we are even able to optimize the result. This upper bound gives an automatic lower bound for the minimal eigenvalue and an upper bound for the negativity as well, see inequality (5.7) and its practical calculation in (5.19). In Subsections 5.2 and 5.3 we show this method in case of a Gaussian kernel, and a Gaussian kernel multiplied by a quadratic polynomial, respectively.
In Subsection 6.1 we define $q_{n,c}$ using the above defined quantities $e_{k}$ and $c$. We prove (i) in Theorem 6.2. This means that the sequence $q_{n,c}$ is monotone increasing and converges to $\lambda_{\min}$ no matter what the value of $c$ is, getting better and better, rigorous lower bounds during the process. We show (ii) with precise lower and upper bounds in Theorem 6.3, implying basically that $q_{n,c}$ depends on $c$ in a linear way if $O$ is positive semidefinite. For (iii) see Theorem 6.4, where we prove a super-exponential speed of convergence
$|q_{n,c}-\lambda_{\min}|\leq\exp\left[-\frac{|\lambda_{\min}|}{c}n\log(n)+\mathcal{O}(n)\right]$
if $O$ is not positive semidefinite. We also calculate and compare our estimates $q_{n,0}$ and $q_{n,c}$ for a concrete operator in Figure 3, obtaining that the sequence $q_{n,0}$ converges extremely fast in practice.
2. Preliminaries
Let $\mathcal{H}$ be a separable, complex Hilbert space [^3] We use the mathematical convention $\langle\lambda f,g\rangle=\lambda\langle f,g\rangle$ and $\langle f,\lambda g\rangle=\lambda^{*}\langle f,g\rangle$, where $\lambda^{*}$ denotes the conjugate of $\lambda$. and let $O\colon\mathcal{H}\to\mathcal{H}$ be a bounded linear operator. We say that $\lambda\in\mathbb{C}$ is an eigenvalue of $O$ if there exists a nonzero $f\in\mathcal{H}$ such that $Of=\lambda f$; then $f$ is an eigenvector corresponding to $\lambda$. The set of eigenvectors
$E_{\lambda}=\{f\in\mathcal{H}:Of=\lambda f\}$
[…]
$\operatorname{Tr}\{O\}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i\geq 0}\langle Of_{i},f_{i}\rangle,$
where the sum is absolutely convergent and independent of the orthonormal basis $f_{i}$, see [deGosson_book, Definition 9.3]. Now assume that $O$ is self-adjoint and trace-class. Every trace-class operator is compact by [deGosson_book, Proposition 9.7]. As $O$ is compact and self-adjoint, the spectral theorem [lax2002functional, Theorem 3 in Chapter 28] implies that it has countably many eigenvalues [^6] Note that in the enumeration $\{\lambda_{i}\}_{i\geq 0}$ we list every eigenvalue $\lambda$ with multiplicity $\dim E_{\lambda}$. $\{\lambda_{i}\}_{i\geq 0}$, and there exists an orthonormal basis consisting of eigenvectors. Choosing such a basis yields
$\operatorname{Tr}\{O\}=\sum_{i\geq 0}\lambda_{i}.$
The main example in this paper is $\mathcal{H}=L^{2}(\mathbb{R}^{d})$, that is, the Hilbert space of the complex-valued square integrable functions defined on $\mathbb{R}^{d}$ endowed with the scalar product $\langle f,g\rangle=\int_{\mathbb{R}^{d}}f(\mathbf{x})g^{*}(\mathbf{x})\,\mathrm{d}\mathbf{x}$, where $z^{*}$ denotes the complex conjugate of $z$. A kernel $K\in L^{2}(\mathbb{R}^{2d})$ defines an integral operator $\widehat{K}\colon L^{2}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$ by the formula
$\big{(}\widehat{K}f\big{)}(\mathbf{x})=\int_{\mathbb{R}^{d}}\,K(\mathbf{x},\mathbf{y})f(\mathbf{y})\,\mathrm{d}\mathbf{y}.$
The operator $\widehat{K}$ is always Hilbert–Schmidt, so compact, see e. g. [deGosson_book]. Thus $\widehat{K}$ has countably many eigenvalues $\{\lambda_{i}\}_{i\geq 0}$, and each eigenvalue $\lambda$ satisfies a Fredholm integral equation
$\int_{\mathbb{R}^{d}}\,K(\mathbf{x},\mathbf{y})f(\mathbf{y})\,\mathrm{d}\mathbf{y}=\lambda f(\mathbf{x})$
with some nonzero functions $f\in L^{2}(\mathbb{R}^{d})$. If the kernel $K$ is continuous, then $\widehat{K}$ is self-adjoint if and only if $K(\mathbf{y},\mathbf{x})=K^{*}(\mathbf{x},\mathbf{y})$ for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. It is a trickier question when $\widehat{K}$ becomes trace-class; somewhat surprisingly adding the natural-looking condition $K\in L^{1}(\mathbb{R}^{2d})$ is not sufficient, see [deGosson_book, Subsection 9.2.2]. The Schwartz space $\mathcal{S}(\mathbb{R}^{d})$ is the set of smooth functions $f\colon\mathbb{R}^{d}\to\mathbb{C}$ with rapidly decreasing mixed partial derivatives, see e. g. [SR, Section V.3]. If $K\in\mathcal{S}(\mathbb{R}^{2d})$, then $\widehat{K}$ is trace-class, see [Gosson_Harmonic, Proposition 287] or [Brislawn, Proposition 1.1] with the remark afterwards. If $K$ is continuous and $\widehat{K}$ is trace-class, then [Duflo, Brislawn] yield the useful formula
$\operatorname{Tr}\big{\{}\widehat{K}\big{\}}=\int_{\mathbb{R}^{d}}K(\mathbf{x},\mathbf{x})\,\mathrm{d}\mathbf{x}.$
Note that $\widehat{K}$ is positive semidefinite if and only if all of its eigenvalues are greater than or equal to zero, which is equivalent to
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}K(\mathbf{x},\mathbf{y})f(\mathbf{x})f^{*}(\mathbf{y})\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}\geq 0\quad\text{for all }f\in L^{2}(\mathbb{R}^{d}).$
For more on Hilbert–Schmidt and trace-class operators see e. g. [Gosson_Harmonic, deGosson_book, Stein].
We say that $\widehat{\rho}$ is a density operator if it is positive semidefinite with $\operatorname{Tr}\big{\{}\widehat{\rho}\big{\}}=1$, that is, we have $\lambda_{i}\geq 0$ and $\sum_{i\geq 0}\lambda_{i}=1$.
Definition 2.1.
A Gaussian kernel $K_{G}\in L^{2}(\mathbb{R}^{2d})$ is of the form
$K_{G}(\mathbf{x},\mathbf{y})=\exp\left\{-\mathbf{r}^{\top}\mathbf{M}\mathbf{r}-\mathbf{V}^{\top}\mathbf{r}+F\right\},$
where $\mathbf{r}=(\mathbf{x},\mathbf{y})^{\top}$, and $\mathbf{M}=\mathbf{M}_{1}+i\mathbf{M}_{2}\in\mathbb{C}^{2d\times 2d}$ is a symmetric complex matrix such that $\mathbf{M}_{1},\mathbf{M}_{2}\in\mathbb{R}^{2d\times 2d}$ and $\mathbf{M}_{1}$ is positive definite, $\mathbf{V}\in\mathbb{C}^{2d}$ is a complex vector, and $F\in\mathbb{C}$ is a complex number. In particular, a self-adjoint Gaussian kernel $K_{G}\in L^{2}(\mathbb{R}^{2d})$ is of the form
$K_{G}(\mathbf{x},\mathbf{y})=$ $\exp\left\{-(\mathbf{x}-\mathbf{y})^{\top}\mathbf{A}(\mathbf{x}-\mathbf{y})-i(\mathbf{x}-\mathbf{y})^{\top}\mathbf{B}(\mathbf{x}+\mathbf{y})\right.$
$\quad\quad\left.-(\mathbf{x}+\mathbf{y})^{\top}\mathbf{C}(\mathbf{x}+\mathbf{y})-i\mathbf{D}^{\top}(\mathbf{x}-\mathbf{y})-\mathbf{E}^{\top}(\mathbf{x}+\mathbf{y})+F\right\},$
where $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{d\times d}$ are real matrices such that $\mathbf{A},\mathbf{C}$ are symmetric and positive definite, $\mathbf{E},\mathbf{D}\in\mathbb{R}^{d}$, and $F\in\mathbb{R}$. We call $K\in L^{2}(\mathbb{R}^{2d})$ a polynomial Gaussian kernel if
(2.1) $K(\mathbf{x},\mathbf{y})=P(\mathbf{x},\mathbf{y})K_{G}(\mathbf{x},\mathbf{y}),$
where $P$ is a self-adjoint polynomial in $2d$ variables, and $K_{G}\in L^{2}(\mathbb{R}^{2d})$ is a self-adjoint Gaussian kernel.
3. Approximating the spectrum
Recall that we use the following notation throughout the paper. Let $O\colon\mathcal{H}\to\mathcal{H}$ be a self-adjoint, trace-class linear operator acting on a separable, complex Hilbert space $\mathcal{H}$. Then $O$ has only countably many real eigenvalues, let us enumerate them with multiplicities as $\{\lambda_{i}\}_{i\geq 0}$. As $O$ is trace-class, we have $\sum_{i\geq 0}|\lambda_{i}|<\infty$. If there are only finitely many eigenvalues $\lambda_{0},\dots,\lambda_{m}$ then let $\lambda_{i}=0$ for each integer $i>m$.
3.1. The quantities $e_{k}$ and their calculation
Following [Newton] let $e_{0}=1$ and for positive integers $k\geq 1$ define
$e_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{0\leq i_{1}<\dots<i_{k}}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$
It was proved in [Newton, Proposition 2.1] that the quantities $e_{k}$ are intimately connected to the positivity of the operator $O$.
Theorem 3.1 (Homa–Balka–Bernád–Károly–Csordás).
The operator $O$ is positive semidefinite if and only if $e_{k}\geq 0$ for all integers $k\geq 1$.
Let $K\in L^{2}(\mathbb{R}^{2d})$ be a self-adjoint kernel, it was showed in [Newton] that the values $e_{k}$ corresponding to the operator $\widehat{K}$ can be calculated by
$e_{k}=\frac{1}{k!}\left|\begin{array}[]{ccccc}M_{1}&1&0&\cdots\\ M_{2}&M_{1}&2&0&\cdots\\ \vdots&&\ddots&\ddots\\ M_{k-1}&M_{k-2}&\cdots&M_{1}&k-1\\ M_{k}&M_{k-1}&\cdots&M_{2}&M_{1}\end{array}\right|,$
where the moments $\{M_{\ell}\}_{1\leq\ell\leq k}$ are defined as
$M_{\ell}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=0}^{\infty}\lambda_{i}^{\ell}=\operatorname{Tr}\big{\{}\widehat{K}^{\ell}\big{\}}=\int_{\mathbb{R}^{\ell d}}\,\left[K(\mathbf{x}_{\ell},\mathbf{x}_{1})\prod_{i=1}^{\ell-1}K(\mathbf{x}_{i},\mathbf{x}_{i+1})\right]\,\mathrm{d}\mathbf{x}_{1}\cdots\mathrm{d}\mathbf{x}_{\ell}.$
Also, it was demonstrated in [Newton] that this calculation is effective for polynomial Gaussian kernels according to Definition 2.1.
Define $g\colon\mathbb{R}\to\mathbb{R}$ as
$g(x)\stackrel{{\scriptstyle\text{def}}}{{=}}\prod_{i=0}^{\infty}(1+\lambda_{i}x).$
By [Simon, Lemma 3.3] we obtain that $g(x)$ is finite for all $x\in\mathbb{R}$ and
(3.1) $g(x)=\sum_{k=0}^{\infty}e_{k}x^{k},$
see also [Newton] for an elementary proof. Taking the algebraic expansion of $\left(\sum_{i=0}^{\infty}|\lambda_{i}|\right)^{k}$ yields that the sequence $e_{k}$ rapidly converges to 00. More precisely, we obtain the following [Simon, Lemma 3.3 (3.4)]:
(3.2) $|e_{k}|\leq\frac{\left(\sum_{i=0}^{\infty}|\lambda_{i}|\right)^{k}}{k!}\quad\textrm{for all }k\geq 1.$
Definition 3.2.
For each integer $n\geq 1$ define the set $\Lambda_{n}\subset\mathbb{R}$ as
$\Lambda_{n}=\left\{-\frac{1}{x}:x\in\mathbb{R}\text{ and }g_{n}(x)=0\right\},\quad\text{where }g_{n}(x)=\sum_{k=0}^{n}e_{k}x^{k}.$
The main idea of this section is that $\Lambda_{n}$ will approximate the closure of the set of non-zero eigenvalues given by
$\left\{-\frac{1}{x}:x\in\mathbb{R}\text{ and }g(x)=\sum_{k=0}^{\infty}e_{k}x^{k}=0\right\}.$
The connection between the two sets is that the polynomial $g_{n}$ is a truncation of the power series $g$. The more precise statement is based on the concept of Hausdorff metric, see Subsection 3.4 for more details.
3.2. A Gaussian example
First we apply our method for a Gaussian operator, where our approximation $\Lambda_{n}$ can be directly compared to the exact spectrum. A self-adjoint Gaussian kernel $K_{G}\colon\mathbb{R}^{2}\to\mathbb{C}$ with trace $1$ is of the form
(3.3) $\begin{split}K_{G}(x,y)=N_{0}\exp&\left\{-A(x-y)^{2}-iB(x^{2}-y^{2})-C(x+y)^{2}\right.\\ &\leavevmode\nobreak\ \left.-iD(x-y)-E(x+y)\right\},\end{split}$
with real parameters $A>0$, $C>0$, $B$, $D$, $E$, where $N_{0}=2\sqrt{\frac{C}{\pi}}\exp\left[-\frac{E^{2}}{4C}\right]$ is a normalizing factor ensuring that $\operatorname{Tr}\big{\{}\widehat{K}_{G}\big{\}}$=1. Let $\beta=\frac{\sqrt{A}-\sqrt{C}}{\sqrt{A}+\sqrt{C}}$, then the eigenvalues are $\lambda_{n}=(1-\beta)\beta^{n}$ where $n\geq 0$, and the quantities $e_{k}$ can be calculated as
$e_{k}=\frac{(1-\beta)^{k}\beta^{\frac{k(k-1)}{2}}}{\prod_{i=1}^{k}(1-\beta^{i})}.$
In Table 1 we considered the parameters $A=1$, $C=4$, $B=D=E=0$.
$k$ $\lambda_{k}$ $\lambda^{(10)}_{k}$ $r_{k}$
00 $1.333333333$ $1.333333333$ 00
$1$ $-0.4444444444$ $-0.4444444444$ 00
$2$ $0.1481481481$ $0.1481481481$ 00
$3$ $-0.04938271605$ $-0.04938271605$ 00
$4$ $0.01646090535$ $0.01646090536$ $6.075\cdot 10^{-10}$
$5$ $-0.00548696845$ $-0.00548696824$ $3.845\cdot 10^{-8}$
$6$ $0.001828989483$ $0.00182897225$ $9.4227\cdot 10^{-6}$
$7$ $-0.00060966316$ $-0.00061011957$ $0.00075$
$8$ $0.00020322$ $0.0002074$ $0.02041$
$9$ $-0.00006774$ $-0.00005448$ $0.19576$
Table 1. The column $\lambda_{k}$ contains the first $10$ eigenvalues ordered from the largest to the smallest absolute value. The column $\lambda^{(10)}_{k}$ contains the elements of the approximation $\Lambda_{10}$ ordered similarly. The last column lists the relative errors $r_{k}=\frac{|\lambda_{k}-\lambda^{(10)}_{k}|}{|\lambda_{k}|}$. Note that, for example, the difference of the real $1$-norm and the one calculated from $\Lambda_{10}$ is just $\sum^{\infty}_{k=0}|\lambda_{k}|-\sum^{9}_{k=0}|\lambda^{(10)}_{k}|=0.000042545$.
3.3. A non-Gaussian example
Now we will test our estimate in case of an integral operator whose kernel is a quadratic polynomial multiplied by a Gaussian. Recall that a polynomial Gaussian kernel $K\colon\mathbb{R}^{2}\to\mathbb{C}$ is given by
$K(x,y)=P(x,y)K_{G}(x,y),$
where $K_{G}(x,y)$ is a self-adjoint Gaussian kernel with trace $1$ of the form (3.3) with real parameters $A,B,C,D,E$ such that $A,C>0$; in our example $P(x,y)$ is a self-adjoint quadratic polynomial given by
$P(x,y)=\frac{1}{N}$ $\left(A_{P}(x-y)^{2}+iB_{P}(x^{2}-y^{2})+C_{P}(x+y)^{2}\right.$
$\,\left.+iD_{P}(x-y)+E_{P}(x+y)+F_{P}\right),$
with real parameters $A_{P},B_{P},C_{P},D_{P},E_{P},F_{P}$ and normalizing factor
$N=F_{P}+\frac{C_{P}-E_{P}E}{2C}+\frac{C_{P}E^{2}}{4C^{2}}.$
We considered the following example. Choose the Gaussian parameters as $A=\frac{3}{2}$, $C=1$, $B=D=E=0$, and let the polynomial parameters be $A_{P}=-1$, $C_{P}=5$, $F_{P}=1$, and $B_{P}=D_{P}=E_{P}=0$. The quantities $e_{k}$ were calculated with Mathematica according to Subsection 3.1. Table 2 summarizes the values $e_{k}$ for $k\in\{0,\dots,10\}$, and $\lambda^{(10)}_{k}$ denotes the approximated eigenvalues in $\Lambda_{10}$.
3.4. A theoretical result
The main goal of this subsection is to show that the set $\Lambda_{n}$ defined in Definition 3.2 converges to the closure of the set of non-zero eigenvalues. Indeed, as the quantities $e_{k}$ from Subsection 3.1 carry no information about the zero eigenvalues, we need to implement a mild modification to the set of eigenvalues. Let $\Lambda\subset\mathbb{R}$ denote the closure of the set of non-zero $\lambda_{i}$, that is,
$\Lambda=\overline{\{\lambda_{i}:i\geq 0\}\setminus\{0\}}.$
This means that $0\in\Lambda$ if and only if there are infinitely many non-zero $\lambda_{i}$.
Definition 3.3.
For two non-empty compact sets $K_{1},K_{2}\subset\mathbb{R}$ let us define their Hausdorff distance as
$\operatorname{d_{H}}(K_{1},K_{2})\stackrel{{\scriptstyle\text{def}}}{{=}}\inf\{r:K_{1}\subset B(K_{2},r)\text{ and }K_{2}\subset B(K_{1},r)\},$
where
$B(A,r)=\{x\in\mathbb{R}:\exists y\in A\text{ such that }|x-y|<r\}.$
Then the non-empty compact subsets of $\mathbb{R}$ endowed with the Hausdorff metric $\operatorname{d_{H}}$ form a complete, separable metric space, see [kechris] for more information.
$k$ $e_{k}$ $\lambda^{(10)}_{k}$
00 $1$ $-0.003010084661$
$1$ $1$ $0.5746896845$
$2$ $0.2493580000$ $0.4133937042$
$3$ $0.002841083333$ $0.008697675828$
$4$ $2.291012500\cdot 10^{-6}$ $0.006056431139$
$5$ $-3.7346411667\cdot 10^{-8}$ $0.0001438482769$
$6$ $-6.493833333\cdot 10^{-12}$ $0.00002652172242$
$7$ $-1.579632937\cdot 10^{-16}$ $1.918870466\cdot 10^{-6}$
$8$ $-3.226636905\cdot 10^{-22}$ $2.677461235\cdot 10^{-7}$
$9$ $-8.414296737\cdot 10^{-29}$ $3.245680477\cdot 10^{-8}$
$10$ $-2.396497465\cdot 10^{-36}$
Table 2. The table contains the first $11$ values of $e_{k}$ and the elements of the tenth approximation $\Lambda_{10}$.
The main goal of this subsection is to prove the following theorem.
Theorem 3.4.
Assume that each non-zero $\lambda_{i}$ has multiplicity $1$. Then $\Lambda_{n}\to\Lambda$ in Hausdorff metric. [^7] This means that, in particular, for large $n$ the set $\Lambda_{n}$ is non-empty.
Proof.
First assume that there are only finitely many non-zero $\lambda_{i}$, say $\lambda_{0},\dots,\lambda_{k-1}$. Then clearly $g_{n}(x)=g(x)$ for all $n\geq k$ for each $x$. Hence $\Lambda_{n}=\Lambda$ for all $n\geq k$.
Now assume that there are infinitely many non-zero $\lambda_{i}$. Fix an arbitrary $\varepsilon>0$, we need to prove that there exists $N\in\mathbb{N}$ such that $\operatorname{d_{H}}(\Lambda_{n},\Lambda)\leq\varepsilon$ for all $n\geq N$. Define
$K_{\varepsilon}=\left\{-\frac{1}{x}:x\in\mathbb{R}\setminus B(\Lambda,\varepsilon)\right\}.$
Clearly $K_{\varepsilon}$ is closed, and $0\in\Lambda$ implies that $K_{\varepsilon}\subset[-1/\varepsilon,1/\varepsilon]$, so $K_{\varepsilon}$ is compact. As $g$ is continuous and non-zero on $K_{\varepsilon}$, we can define $\theta\stackrel{{\scriptstyle\text{def}}}{{=}}\min\{|g(x)|:x\in K_{\varepsilon}\}>0$. Since $g_{n}\to g$ uniformly on $K_{\varepsilon}$, there exists $N_{1}\in\mathbb{N}$ such that $|g_{n}(x)-g(x)|<\theta$ for all $n\geq N_{1}$ and $x\in K_{\varepsilon}$. Therefore, the definition of $\theta$ yields $K_{\varepsilon}\cap\{x:g_{n}(x)=0\}=\emptyset$ for all $n\geq N_{1}$. Hence the definitions of $K_{\varepsilon}$ and $\Lambda_{n}$ imply for each $n\geq N_{1}$ that
(3.4) $\Lambda_{n}\subset B(\Lambda,\varepsilon).$
Now choose finitely many non-zero $\lambda_{i}$ (we may assume that $\lambda_{0},\dots,\lambda_{k-1}$) and $\delta>0$ such that
(3.5) $[\lambda_{i}-\delta,\lambda_{i}+\delta]\cap\Lambda=\{\lambda_{i}\}\quad\text{for all }0\leq i<k,$
and for any points $x_{i}\in[\lambda_{i}-\delta,\lambda_{i}+\delta]$ we have
(3.6) $\Lambda\subset B\left(\cup_{i=0}^{k-1}\{x_{i}\},\varepsilon\right).$
Define a $2k$ element set
$C\stackrel{{\scriptstyle\text{def}}}{{=}}\left\{\frac{1}{-\lambda_{i}\pm\delta}:0\leq i<k\right\},$
and let $I_{i}$ be the open interval with endpoints $1/(-\lambda_{i}\pm\delta)$ for any $0\leq i<k$. As $\lambda_{i}$ has multiplicity $1$, (3.5) yields that $g(\frac{1}{-\lambda_{i}-\delta})$ and $g(\frac{1}{-\lambda_{i}+\delta})$ are non-zero and have opposite signs for all $i$. As $g_{n}\to g$ uniformly on $C$, we infer that there exists $N_{2}\in\mathbb{N}$ such that for each $n\geq N_{2}$ the values $g_{n}(\frac{1}{-\lambda_{i}-\delta})$ and $g_{n}(\frac{1}{-\lambda_{i}+\delta})$ are non-zero and have opposite signs for all $0\leq i<k$. Thus for any $n\geq N_{2}$ for all $0\leq i<k$ by the continuity of $g_{n}$ we can define $z_{i,n}\in I_{i}$ such that $g_{n}(z_{i,n})=0$. Let $x_{i,n}=-1/z_{i,n}$, then clearly $\{x_{i,n}:0\leq i<k\}\subset\Lambda_{n}$, and $x_{i,n}\in(\lambda_{i}-\delta,\lambda_{i}+\delta)$ for all $0\leq i<k$. Then (3.6) implies that
(3.7) $\Lambda\subset B(\{x_{i,n}:0\leq i<k\},\varepsilon)\subset B(\Lambda_{n},\varepsilon).$
Finally, let $N=\max\{N_{1},N_{2}\}$, then (3.4) and (3.7) imply for all $n\geq N$ that
$\operatorname{d_{H}}(\Lambda_{n},\Lambda)\leq\varepsilon.$
As $\varepsilon>0$ was arbitrary, the proof is complete. ∎
4. A non-monotone approximation of the minimal eigenvalue
We apply the method of Section 3 to estimate the minimal eigenvalue. Let the operator $O$ and its real eigenvalues $\lambda_{i}$ be the same as in Section 3. Recall that $\lambda_{\min}=\min\{0,\lambda_{i}:i\geq 0\}$, and note that
(4.1) $\lambda_{\min}=\frac{-1}{\min\{x>0:g(x)=0\}},$
where we use the convention $\min\emptyset=+\infty$. Let us recall the following definition.
Definition 4.1.
For all integers $n\geq 1$ define
(4.2) $q_{n,0}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{-1}{\min\left\{x>0:g_{n}(x)=0\right\}},\quad\text{where }g_{n}(x)=\sum_{k=0}^{n}e_{k}x^{k}.$
We consider the approximation of $\lambda_{\min}$ by $q_{n,0}$. In the next theorem we prove that $q_{n,0}\to\lambda_{\min}$ with super-exponential speed under the slight technical condition that $\lambda_{\min}<0$ has multiplicity one. First we need the following easy fact.
Fact 4.2.
Let $n\geq 0$ be an integer and let $0\leq x\leq\frac{n+2}{2}$. Then
$\sum_{k=n+1}^{\infty}\frac{x^{k}}{k!}\leq\frac{2x^{n+1}}{(n+1)!}$
Proof.
We can use a geometric sum for the estimate as
$\sum_{k=n+1}^{\infty}\frac{x^{k}}{k!}\leq\frac{x^{n+1}}{(n+1)!}\sum_{i=0}^{\infty}\left(\frac{x}{n+2}\right)^{i}\leq\frac{2x^{n+1}}{(n+1)!}.\qed$
Theorem 4.3.
Assume that $\lambda_{\min}<0$ has multiplicity one. Then $q_{n,0}\to\lambda_{\min}$. Moreover,
(4.3) $|q_{n,0}-\lambda_{\min}|\leq\exp[-n\log(n)+\mathcal{O}(n)].$
Proof.
First we prove that $q_{n,0}\to\lambda_{\min}$. Assume that $\lambda_{0}=\lambda_{\min}$ and fix $\gamma_{0}$ such that $\lambda_{0}<\gamma_{0}<\min\{0,\lambda_{i}:i\geq 1\}$. Define
$\theta_{0}=-\frac{1}{\lambda_{0}}\quad\text{and}\quad\theta_{1}=-\frac{1}{\gamma_{0}}.$
Since $g$ has no root between $\theta_{0}$ and $\theta_{1}$, and $\theta_{0}$ is the smallest root of $g$ with multiplicity $1$, we obtain that $g(x)>0$ for all $0\leq x<\theta_{0}$ and $g(x)<0$ for all $\theta_{0}<x\leq\theta_{1}$. Fix an arbitrary $0<\varepsilon<\min\{\theta_{0},\theta_{1}-\theta_{0}\}$, and define $\delta>0$ as
$\delta\stackrel{{\scriptstyle\text{def}}}{{=}}\min\{|g(x)|:x\in[0,\theta_{0}-\varepsilon]\cup\{\theta_{0}+\varepsilon\}\}.$
Since $g_{n}$ uniformly converges to $g$ on the interval $[0,\theta_{0}+\varepsilon]$, we obtain that there exists an integer $N=N(\varepsilon)\geq 0$ such that
(4.4) $|g_{n}(x)-g(x)|<\delta\quad\text{for all }x\in[0,\theta_{0}+\varepsilon]\text{and }n\geq N.$
The definition of $\delta$ and (4.4) imply that for all $n\geq N$ we have $g_{n}(x)>0$ for all $x\in[0,\theta_{0}-\varepsilon]$ and $g_{n}(\theta_{0}+\varepsilon)<0$. Hence $\min\left\{x>0:g_{n}(x)=0\right\}\in[\theta_{0}-\varepsilon,\theta_{0}+\varepsilon]$ for all $n\geq N$. Taking the limit $\varepsilon\to 0+$ yields that $\min\left\{x>0:g_{n}(x)=0\right\}\to\theta_{0}$ as $n\to\infty$, which implies that $q_{n,0}\to\lambda_{0}$ by definition.
Now we prove the upper bound (4.3). Let $F(x)=|\lambda_{0}|\prod_{i\geq 1}(1+\lambda_{i}x)$, and let
$\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\min\left\{F(x):0\leq x\leq\theta_{1}\right\}.$
As $F$ is continuous and positive on the interval $[0,\theta_{1}]$, we obtain that $\Delta>0$. Consider $x=\theta_{0}+s$, where $s\in[-\theta_{0},\theta_{1}-\theta_{0}]$. Then the definition of $\Delta$ yields
(4.5) $|g(x)|=\left|\prod_{i=0}^{\infty}(1+\lambda_{i}x)\right|=|s\lambda_{0}|\prod_{i\geq 1}(1+\lambda_{i}x)\geq\Delta|s|.$
Define
$L\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=0}^{\infty}|\lambda_{i}|,$
and assume that $n>2\theta_{1}L$. Then $Lx\leq L\theta_{1}<\frac{n}{2}$ for every $x\in[0,\theta_{1}]$. Thus (3.2), Fact 4.2, and the inequality $n!\geq e\left(\frac{n}{e}\right)^{n}$ imply that for all $0\leq x\leq\theta_{1}$ we have
(4.6) $|g_{n}(x)-g(x)|=\left|\sum_{k=n+1}^{\infty}e_{k}x^{k}\right|\leq\sum_{k=n+1}^{\infty}\frac{(Lx)^{k}}{k!}\leq 2\frac{(Lx)^{n+1}}{(n+1)!}<\left(\frac{eL\theta_{1}}{n+1}\right)^{n+1}.$
Now define $\varepsilon_{n}>0$ as
(4.7) $\varepsilon_{n}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{\Delta}\left(\frac{eL\theta_{1}}{n+1}\right)^{n+1}.$
Then (4.5), (4.6), and (4.7) yield that for all large enough $n$ we have $g_{n}(x)>0$ for all $0\leq x\leq\theta_{0}-\varepsilon_{n}$ and $g_{n}(\theta_{0}+\varepsilon_{n})<0$; so we can define
$r_{n}\stackrel{{\scriptstyle\text{def}}}{{=}}\min\{x>0:g_{n}(x)=0\}\in[\theta_{0}-\varepsilon_{n},\theta_{0}+\varepsilon_{n}].$
Then $r_{n}\in[\theta_{0}-\varepsilon_{n},\theta_{0}+\varepsilon_{n}]$, and the definition of $\varepsilon_{n}$ with some straightforward analysis imply that for all large enough $n$ we have
$|q_{n,0}-\lambda_{0}|=\left|\frac{1}{\theta_{0}}-\frac{1}{r_{n}}\right|\leq\left|\frac{1}{\theta_{0}}-\frac{1}{\theta_{0}-\varepsilon_{n}}\right|\leq 2|\lambda_{0}|^{2}|\varepsilon_{n}|=\exp[-n\log(n)+\mathcal{O}(n)].$
The proof of the theorem is complete. ∎
Figure 3 will compare this approximation with the one in Section 6.
Remark 4.4.
Note that the proof of an analogous theorem would also work if we only assumed that $\lambda_{\min}<0$ has odd multiplicity $m$ in the sequence $\{\lambda_{i}\}_{i\geq 0}$, with the only difference that $n\log(n)$ need to be replaced by $\frac{1}{m}n\log(n)$ in (4.3). However, having multiplicity bigger than one does not seem to be a natural condition for the minimal eigenvalue of a linear operator.
5. Approximation of the Schatten–von Neumann 1-norm
The main goal of this section is to give a method which can provide rigorous upper bounds for the Schatten–von Neumann $1$-norm, and to demonstrate that it works effectively for polynomial Gaussian operators, see Definition 2.1. We use this section as a preparation for Section 6, but it can be interesting in its own right. Let $\mathcal{H}$ be a separable, complex Hilbert space. For the following definition see for example [SR, Chapter VI].
Definition 5.1.
Assume that $O\colon\mathcal{H}\to\mathcal{H}$ is a Hilbert–Schmidt operator. Let $\{s_{i}\}_{i\geq 0}$ be the singular values of $O$, that is, the sequence of the square roots of the eigenvalues of $OO^{\dagger}$ with multiplicities. Then we can define the finite $2$-norm of $O$ as
$\|O\|_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}\Bigg{(}\sum_{i\geq 0}s_{i}^{2}\Bigg{)}^{\frac{1}{2}}.$
If $O$ is trace-class, we can define its Schatten–von Neumann $1$-norm as
$\|O\|_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i\geq 0}s_{i}.$
By [SR, Theorem VI.23] for $K\in L^{2}(\mathbb{R}^{2d})$ the $2$-norm of the Hilbert–Schmidt integral operator $\widehat{K}$ can be calculated as
(5.1) $\big{\|}\widehat{K}\big{\|}^{2}_{2}=\int_{\mathbb{R}^{2d}}|K(\mathbf{x},\mathbf{y})|^{2}\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}.$
Now assume that $O$ is a self-adjoint, trace-class operator, then clearly we obtain that $\|O\|_{1}=\sum_{i\geq 0}|\lambda_{i}|<\infty$, where $\{\lambda_{i}\}_{i\geq 0}$ are the eigenvalues with multiplicities. The main goal of this section is to find an upper bound for $\|O\|_{1}$ which can be effectively calculated. By [SR, Theorem VI. 22] the operator $O$ can be written as a product
(5.2) $O=O_{1}O_{2},$
where $O_{1},O_{2}\colon\mathcal{H}\to\mathcal{H}$ are Hilbert–Schmidt operators, so $\|O_{i}\|_{2}<\infty$ for $i=1,2$. Hölder’s inequality [SR, Problem 28 in Chapter VI] yields
(5.3) $\|O\|_{1}\leq\|O_{1}\|_{2}\|O_{2}\|_{2},$
giving an upper bound for the Schatten–von Neumann $1$-norm of $O$. However, it is a hard problem to find a decomposition (5.2) where we can also calculate or estimate the norms $\|O_{1}\|_{2}$ and $\|O_{2}\|_{2}$.
We propose a solution to self-adjoint trace-class integral operators $\widehat{K}$ with kernels consisting of a polynomial multiplied by a Gaussian kernel, see Subsection 5.1. Then a decomposition to Hilbert–Schmidt operators
(5.4) $\widehat{K}=\widehat{K}_{1}\widehat{K}_{2}$
implies that the kernels satisfy
(5.5) $K(\mathbf{x},\mathbf{y})=\int_{\mathbb{R}^{d}}K_{1}(\mathbf{x},\mathbf{z})K_{2}(\mathbf{z},\mathbf{y})\,\mathrm{d}\mathbf{z}.$
Applying Hölder’s inequality (5.3) to the decomposition (5.4), and using the formula (5.1) for $K_{1}$ and $K_{2}$ yield
(5.6) $\big{\|}\widehat{K}\big{\|}_{1}\leq\left(\int_{\mathbb{R}^{2d}}|K_{1}(\mathbf{x},\mathbf{y})|^{2}\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{2d}}|K_{2}(\mathbf{x},\mathbf{y})|^{2}\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}\right)^{\frac{1}{2}}.$
This upper bound has a direct connection to the minimal eigenvalue problem. We obtain an immediate lower bound for $\lambda_{\min}$ and upper bound for the negativity $\mathcal{N}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=0}^{\infty}|\min\{\lambda_{i},0\}|$ as
(5.7) $|\lambda_{\min}|\leq\mathcal{N}=\frac{\big{\|}\widehat{K}\big{\|}_{1}-\operatorname{Tr}\big{\{}\widehat{K}\big{\}}}{2}\leq\frac{\big{\|}\widehat{K}_{1}\big{\|}_{2}\big{\|}\widehat{K}_{2}\big{\|}_{2}-\operatorname{Tr}\big{\{}\widehat{K}\big{\}}}{2}.$
It is interesting that while our operator $\widehat{K}$ is always self-adjoint, often there are only decompositions of the form (5.4) in which neither $\widehat{K}_{1}$ nor $\widehat{K}_{2}$ is self-adjoint. We will see more sophisticated estimates of $\lambda_{\min}$ in Section 6, where we will also use an upper bound for $\big{\|}\widehat{K}\big{\|}_{1}$ coming from (5.4) and (5.6). In the rest of this section we show a method how to calculate such upper bounds.
5.1. Decomposition and optimisation
Given a polynomial Gaussian kernel $K\in L^{2}(\mathbb{R}^{2d})$ of the form (2.1), we want to decompose the operator $\widehat{K}$ to a product according to (5.4). We will search the decomposition in the form
$K_{1}(\mathbf{x},\mathbf{y})=P_{1}(\mathbf{x},\mathbf{y})K_{G_{1}}(\mathbf{x},\mathbf{y})\quad\text{and}\quad K_{2}(\mathbf{x},\mathbf{y})=P_{2}(\mathbf{x},\mathbf{y})K_{G_{2}}(\mathbf{x},\mathbf{y}),$
where $K_{G_{1}},K_{G_{2}}$ are not necessarily self-adjoint Gaussian kernels in $2d$ variables, and $P_{1},P_{2}$ are polynomials in $2d$ variables such that $\deg P_{1}+\deg P_{2}=\deg P$. In particular, if $\deg P_{1}=0$ or $\deg P_{2}=0$ then the decomposition leads us to a system of linear equations which can be solved effectively, and we can even try to minimize $\big{\|}\widehat{K}_{1}\big{\|}_{2}\big{\|}\widehat{K}_{2}\big{\|}_{2}$ coming from the different decompositions of $\widehat{K}$. We demonstrate this algorithm in the $d=1$ case together with a minimization in the following subsections.
5.2. Gaussian operators in dimension one
Recall that a self-adjoint Gaussian kernel $K_{G}\colon\mathbb{R}^{2}\to\mathbb{C}$ is of the form
(5.8) $\begin{split}K_{G}(x,y)=N_{0}\exp&\left\{-A(x-y)^{2}-iB(x^{2}-y^{2})-C(x+y)^{2}\right.\\ &\leavevmode\nobreak\ \left.-iD(x-y)-E(x+y)\right\},\end{split}$
with real parameters $A>0$, $C>0$, $B$, $D$, $E$, where $N_{0}=2\sqrt{\frac{C}{\pi}}\exp\left[-\frac{E^{2}}{4C}\right]$ is a normalizing factor ensuring that $\operatorname{Tr}\big{\{}\widehat{K}_{G}\big{\}}$=1. The eigenvalue equation was solved in [Bernád2018]; the eigenvalues are given by
$\lambda_{i}=\frac{2\sqrt{C}}{\sqrt{A}+\sqrt{C}}\left(\frac{\sqrt{A}-\sqrt{C}}{\sqrt{A}+\sqrt{C}}\right)^{i},\quad i=0,1,\ldots.$
Correspondingly, the Schatten–von Neumann $1$-norm of the Gaussian $\widehat{K}_{G}$ is
(5.9) $\big{\|}\widehat{K}_{G}\big{\|}_{1}=\sum_{i=0}^{\infty}|\lambda_{i}|=\begin{cases}1&\text{if }A\geq C,\\ \sqrt{\frac{C}{A}}&\text{if }C>A.\end{cases}$
By (5.5) we consider the decomposition of $K_{G}$ as
(5.10) $K_{G}(x,y)=\int_{-\infty}^{\infty}\,K_{1}(x,z)K_{2}(z,y)\,\mathrm{d}z.$
We define $K_{1}$ and $K_{2}$ containing the free parameter $w\in\mathbb{R}$ as follows. Let
$K_{1}(x,y)=N_{1}\exp\left(-A_{1}x^{2}-B_{1}y^{2}-C_{1}xy-D_{1}x-E_{1}y\right),$
with
(5.11) $\begin{split}A_{1}&=w+iB+\frac{AC}{w}\\ B_{1}&=w-iB+\frac{AC}{w},\\ C_{1}&=-2\left(w-\frac{AC}{w}\right),\\ D_{1}&=iD+\frac{AE}{w},\\ E_{1}&=-iD+\frac{AE}{w},\\ N_{1}&=\frac{2\sqrt{C}}{\pi}\sqrt{\frac{(w^{2}-AC)^{2}}{w(A-w)(C-w)}}\exp{\left(-\frac{\left(AC-w^{2}\right)E^{2}}{4w(C-w)C}\right)}.\end{split}$
Similarly let
$K_{2}(x,y)=\exp\left(-A_{2}x^{2}-B_{2}y^{2}-C_{2}xy-D_{2}x-E_{2}y\right)$
with
(5.12) $\begin{split}A_{2}&=iB+\frac{(A-w)C}{C-w}+\frac{A(C-w)}{A-w},\\ B_{2}&=-iB+\frac{(A-w)C}{C-w}+\frac{A(C-w)}{A-w},\\ C_{2}&=2\left(\frac{(A-w)C}{C-w}-\frac{A(C-w)}{A-w}\right),\\ D_{2}&=iD+\frac{(A-w)E}{C-w},\\ E_{2}&=-iD+\frac{(A-w)E}{C-w}.\end{split}$
It is easy to check that (5.10) holds and the integral is finite if $w$ satisfies
(5.13) $0<w<\min(A,C)\qquad\text{or}\qquad 0<\max(A,C)<w.$
Straightforward calculation leads to
(5.14) $\begin{split}\iint_{\mathbb{R}^{2}}\left|K_{1}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y=&\frac{C(w^{2}-AC)^{2}}{\pi\sqrt{AC}w(A-w)(C-w)}\exp{\left(-\frac{(A-w)E^{2}}{2(C-w)C}\right)},\\ \iint_{\mathbb{R}^{2}}\left|K_{2}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y=&\frac{\pi}{4\sqrt{AC}}\exp{\left(\frac{(A-w)E^{2}}{2(C-w)C}\right)},\end{split}$
provided that $w$ is taken from the allowed intervals (5.13). By (5.14) we can calculate the product
$\iint_{\mathbb{R}^{2}}\left|K_{1}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y\iint_{\mathbb{R}^{2}}\left|K_{2}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y=\frac{(w^{2}-AC)^{2}}{4Aw(A-w)(C-w)}.$
We can minimize this in $w$ on $[0,\min[A,C]]\cup[\max[A,C],\infty]$, which gives
(5.15) $w_{\min}^{G}=\begin{cases}A\mp\sqrt{A^{2}-AC}&\qquad\text{if }A\geq C,\\ C\mp\sqrt{C^{2}-AC}&\qquad\text{if }C>A.\end{cases}$
The minimum value are the same at both (±)plus-or-minus($\pm$)( ± ) values. Therefore, using Hölder’s inequality (5.6) we obtain the upper bound
$\big{\|}\widehat{K}_{G}\big{\|}_{1}$ $\leq\left.\sqrt{\iint_{\mathbb{R}^{2}}\left|K_{1}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y\iint_{\mathbb{R}^{2}}\left|K_{2}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y}\right|_{w=w_{\min}^{G}}$
=$=\begin{cases}1&\text{if }A\geq C,\\ \sqrt{\frac{C}{A}}&\text{if }C<A.\end{cases}$
Comparing this with (5.9) shows that our upper bound equals to the exact value of the norm $\big{\|}\widehat{K}_{G}\big{\|}_{1}$.
5.3. Quadratic polynomials multiplied by a Gaussian
Our second example is the polynomial Gaussian kernel $K\colon\mathbb{R}^{2}\to\mathbb{C}$ given by
(5.16) $K(x,y)=P(x,y)K_{G}(x,y),$
where $K_{G}(x,y)$ is the self-adjoint Gaussian kernel with trace $1$ from (5.8), and $P(x,y)$ is the self-adjoint polynomial given by
$P(x,y)=\frac{1}{N}$ $\left(A_{P}(x-y)^{2}+iB_{P}(x^{2}-y^{2})+C_{P}(x+y)^{2}\right.$
$\,\left.+iD_{P}(x-y)+E_{P}(x+y)+F_{P}\right),$
with real parameters $A_{P},B_{P},C_{P},D_{P},E_{P},F_{P}$ and normalizing factor
$N=F_{P}+\frac{C_{P}-E_{P}E}{2C}+\frac{C_{P}E^{2}}{4C^{2}}.$
This ensures that $\operatorname{Tr}\big{\{}\widehat{K}\big{\}}=1$. Let us decompose $\widehat{K}=\widehat{K}_{1}\widehat{K}_{2}$, by (5.5) the kernels must satisfy
(5.17) $K(x,y)=\int_{-\infty}^{\infty}K_{1}(x,z)K_{2}(z,y)\,\mathrm{d}z.$
We assume that $K_{1}$ is a Gaussian kernel and $K_{2}$ is a Gaussian multiplied by a quadratic polynomial of the form
$K_{1}(x,y)$ $=N^{\prime}_{1}\exp\left(-A_{1}x^{2}-B_{1}y^{2}-C_{1}xy-D_{1}x-E_{1}y\right);$
$K_{2}(x,y)$ $=P_{2}(x,y)\exp\left(-A_{2}x^{2}-B_{2}y^{2}-C_{2}xy-D_{2}x-E_{2}y\right),$
where
$P_{2}(x,y)=$ $A_{P_{2}}(x-y)^{2}+iB_{P_{2}}(x^{2}-y^{2})+C_{P_{2}}(x+y)^{2}$
$+iD_{P_{2}}(x-y)+E_{P_{2}}(x+y)+F_{P_{2}},$
and all the parameters are real. Choose the exponential parameters $A_{1},\dots,E_{1}$, and $A_{2},\dots,E_{2}$ according to (5.11)-(5.12). By a straightforward calculation the normalizing factor $N^{\prime}_{1}$ and the polynomial parameters $A_{P_{2}},\dots,F_{P_{2}}$ are uniquely determined by the decomposition (5.17): one obtains linear equations for the polynomial parameters. Then the only free parameter is $w$ (as before), which allows minimization. Allowed regions for $w$ are still given by (5.13), where $A$ and $C$ are real parameters of the Gaussian $K_{G}(x,y)$. The quantity
$R(w)=\left(\iint_{\mathbb{R}^{2}}\left|K_{1}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y\right)\left(\iint_{\mathbb{R}^{2}}\left|K_{2}(x,y)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y\right)$
can be easily calculated. It is a lengthy rational function of the free parameter $w$, thus we do not quote it here. Finding the minimum point $w_{\min}$ of $R(w)$ in the allowed regions (5.13) give an optimized upper bound to the norm $\big{\|}\widehat{K}\big{\|}_{1}$. Note that the value of $w_{\min}$ is different from $w_{\min}^{G}$ given in (5.15), our minimum is affected by the polynomial parameters as well. On Figure 1 we display upper bounds of the Schatten–von Neumann $1$-norm for a family of quadratic Gaussian operators of the form (5.16). The decomposition (5.17) is not as optimal as in our first example. It can be proved that at $C_{P}=1$ the the operator $\widehat{K}$ is positive semidefinite, thus the $1$-norm is exactly $1$, but our decomposition (5.17) after the minimization leads to the upper bound
(5.18) $\big{\|}\widehat{K}\big{\|}_{1}\leq 1.04054.$
Then (5.7) yields a lower bound for the minimal eigenvalue $\lambda_{\min}$ and an upper bound for the negativity $\mathcal{N}$ as follows:
(5.19) $|\lambda_{\min}|\leq\mathcal{N}\leq\frac{1.04054-1}{2}=0.02027.$
Figure 1. Upper bounds for the Schatten–von Neumann $1$-norm of the operator $\widehat{K}$ as a function of the polynomial parameter $C_{P}$. The Gaussian parameters are given as $A=\frac{3}{2}$, $C=1$, $B=D=E=0$, and the polynomial parameters are $A_{P}=-1$, $F_{P}=1$, and $B_{P}=D_{P}=E_{P}=0$. Note that the operator $\widehat{K}$ is easily seen to be positive semidefinite at $C_{P}=1$. Here $L_{1}^{(1)}=\sqrt{R(w_{\text{min}})}$ is the solid line, $L_{1}^{(2)}=\sqrt{R(w_{\text{min}}^{G})}$ is the dashed curve. The dash-dotted curve $L_{1}^{(3)}$ shows an estimate obtained from numerical diagonalization of the operator $\widehat{K}$, for the details see Subsection 6.2.
6. A monotone approximation of the minimal eigenvalue
Let $O\colon\mathcal{H}\to\mathcal{H}$ be a self-adjoint, trace-class linear operator acting on a separable, complex Hilbert space $\mathcal{H}$. Then $O$ has countably many real eigenvalues $\{\lambda_{i}\}_{i\geq 0}$ satisfying $\sum_{i\geq 0}|\lambda_{i}|<\infty$. If there are only finitely many eigenvalues $\lambda_{0},\dots,\lambda_{m}$ then define $\lambda_{i}=0$ for each integer $i>m$. From now on let us fix $c>0$ satisfying
(6.1) $\sum_{i=0}^{\infty}|\lambda_{i}|\leq c.$
6.1. The estimate $q_{n,c}$ and its speed of convergence
Definition 6.1.
For $n\geq 0$ define $h_{n}=h_{n,c}\colon[0,\infty)\to\mathbb{R}$ as
$h_{n,c}(x)\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{k=0}^{n}e_{k}x^{k}-\sum_{k=n+1}^{\infty}\frac{(cx)^{k}}{k!},$
and define $q_{n}=q_{n,c}$ as [^8] If it does not cause confusion, we simply use $h_{n}$ and $q_{n}$ instead of $h_{n,c}$ and $q_{n,c}$, respectively.
(6.2) $q_{n,c}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{-1}{\min\{x>0:h_{n,c}(x)=0\}}.$
The following theorem shows that $q_{n}$ is a monotone, asymptotically sharp lower bound of the minimal eigenvalue $\lambda_{\min}$.
Theorem 6.2.
The sequence $q_{n}$ is monotone increasing and
$\lim_{n\to\infty}q_{n}=\lambda_{\min}.$
Proof.
First we show that for all $x\geq 0$ and $n\geq 0$ we have
(6.3) $g(x)\geq h_{n+1}(x)\geq h_{n}(x).$
Let us fix $x\geq 0$ and $n\geq 0$. The definition of $g,h_{n}$ and (3.2) imply that
$g(x)=\sum_{k=0}^{\infty}e_{k}x^{k}\geq\sum_{k=0}^{n}e^{k}x^{k}-\sum_{k=n+1}^{\infty}\frac{(cx)^{k}}{k!}=h_{n}(x).$
Inequality (3.2) also implies that
$h_{n+1}(x)-h_{n}(x)=\left(e_{n+1}+\frac{c^{n+1}}{(n+1)!}\right)x^{n+1}\geq 0,$
thus (6.3) holds. Therefore the definition of $q_{n}$, inequalities (4.1) and (6.3), and $g(0)=h_{n}(0)=1$ imply that $q_{n}$ is monotone increasing and $q_{n}\leq\lambda_{\min}$ for all $n\geq 0$. Finally, $q_{n}\to\lambda_{\min}$ follows from the fact that $h_{n}$ uniformly converges to $g$ on any bounded interval $I\subset[0,\infty)$. ∎
The following theorem states that if $\lambda_{\min}=0$ then $q_{n}\approx-\frac{c}{n}$, see Figure 2 for a numerical example.
Theorem 6.3.
Let $\lambda_{\min}=0$. Then for any integer $n\geq 0$ we have
$\frac{c}{n+1}\leq|q_{n}|\leq\frac{ec}{n+1}.$
Moreover, the lower bound of $|q_{n}|$ holds without the assumption $\lambda_{\min}=0$, too.
Proof.
First we prove the upper bound. Let $\psi_{n}(x)=\sum_{k=n+1}^{\infty}\frac{x^{k}}{k!}$ and let $b_{n}$ be the only positive solution to the equation $\psi_{n}(x)=1$. As $e_{k}\geq 0$ for all $k$, we have $\sum_{k=0}^{n}e_{k}x^{k}\geq 1$ for all $x\geq 0$. Therefore, the definition of $h_{n}$ implies that
$\min\{x>0:h_{n}(x)=0\}\geq\min\{x>0:\psi_{n}(cx)=1\}=\frac{b_{n}}{c},$
which yields
(6.4) $|q_{n}|\leq\frac{c}{b_{n}}.$
Before proving a lower bound for $b_{n}$ we show that
(6.5) $b_{n}\leq\frac{n+2}{2}.$
Let $x_{n}=\frac{n+2}{2}$, we need to prove that $\psi_{n}(x_{n})\geq 1$. By a straightforward estimate
$\psi_{n}(x_{n})\geq\frac{(x_{n})^{n+1}}{(n+1)!}=\frac{(n+2)^{n+1}}{2^{n+1}(n+1)!},$
so it is enough to show that the sequence $a_{n}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{(n+2)^{n+1}}{2^{n+1}(n+1)!}$ satisfies $a_{n}\geq 1$ for all $n$. Indeed, $a_{0}=1$ and using the monotonicity of the sequence $\left(1+\frac{1}{n}\right)^{n}$ for all $n\geq 1$ we obtain
$\frac{a_{n}}{a_{n-1}}=\frac{1}{2}\left(1+\frac{1}{n+1}\right)^{n+1}>1,$
so $a_{n}\geq 1$. Therefore (6.5) holds.
By Fact 4.2 we have
$\psi_{n}(x)\leq\frac{2x^{n+1}}{(n+1)!}\quad\text{for all }0\leq x\leq\frac{n+2}{2}.$
Using this together with (6.5), and the well-known inequality $n!\geq e\left(\frac{n}{e}\right)^{n}$ we obtain
$1=\psi_{n}(b_{n})\leq 2\frac{(b_{n})^{n+1}}{(n+1)!}<\left(\frac{eb_{n}}{n+1}\right)^{n+1}.$
Thus $b_{n}\geq\frac{n+1}{e}$. Hence (6.4) yields
$|q_{n}|\leq\frac{c}{b_{n}}\leq\frac{ec}{n+1},$
which finishes the proof of the upper bound.
Finally, we will prove the lower bound. For integers $n\geq 0$ and reals $x\geq 0$ consider the function
$f_{n}(x)=\sum_{k=0}^{n}\frac{x^{k}}{k!}-\sum_{k=n+1}^{\infty}\frac{x^{k}}{k!},$
and let
$d_{n}=\min\{x>0:f_{n}(x)=0\}.$
By (3.2) we obtain that $h_{n}(x)\leq f_{n}(cx)$. This inequality and $h_{n}(0)=f_{n}(0)=1$ imply that
$\min\{x>0:h_{n}(x)=0\}\leq\min\{x>0:f_{n}(cx)=0\}\leq\frac{d_{n}}{c},$
which yields that
(6.6) $|q_{n}|\geq\frac{c}{d_{n}}.$
Thus it is enough to show that
(6.7) $d_{n}\leq n+1,$
for which it is sufficient to see that $f_{n}(n+1)\leq 0$. In order to see this, it is enough to prove that for all integers $0\leq m\leq n$ we have
(6.8) $\frac{(n+1)^{n-m}}{(n-m)!}\leq\frac{(n+1)^{n+1+m}}{(n+1+m)!},$
since summarizing (6.8) from $m=0$ to $n$ clearly yields
$\sum_{k=0}^{n}\frac{(n+1)^{k}}{k!}\leq\sum_{k=n+1}^{2n+1}\frac{(n+1)^{k}}{k!}<\sum_{k=n+1}^{\infty}\frac{(n+1)^{k}}{k!}.$
Inequality (6.8) is equivalent to
$\prod_{i=-m}^{m}(n+1+i)\leq(n+1)^{2m+1},$
which clearly follows from the identity
$(n+1-i)(n+1+i)=(n+1)^{2}-i^{2}\leq(n+1)^{2}.$
Hence we proved (6.7). Inequalities (6.6) and (6.7) imply that
$|q_{n}|\geq\frac{c}{d_{n}}\geq\frac{c}{n+1},$
so the proof of the lower bound is also complete. ∎
Figure 2. Consider the positive semidefinite polynomial Gaussian operator $\widehat{K}$ of the form (5.16) with parameters $A=\frac{3}{2}$, $C=1$, $B=D=E=0$, $A_{P}=-1$, $C_{P}=1$, $F_{P}=1$, and $B_{P}=D_{P}=E_{P}=0$. By (5.18) the upper bound of $\big{\|}\widehat{K}\big{\|}_{1}$ can be chosen to be $c=1.04054$. We plotted $|q_{n}|=|q_{n,c}|$, $q_{n}^{<}=\frac{c}{n+1}$, and $q_{n}^{>}=\frac{ec}{n+1}$ as a function of $n$ calculated from (6.2).
In the case of $\lambda_{\min}<0$ the following theorem establishes that $q_{n}\to\lambda_{\min}$ with super-exponential speed. On the good side, $q_{n,c}$ is monotone, so our method provides a rigorous lower bound for $\lambda_{\min}$. On the bad side, $q_{n,0}$ approximates $\lambda_{\min}$ much better than $q_{n,c}$, and the main reason of this is that $q_{n,0}$ does not depend on the value of $c$ through the sub-optimal estimate (3.2). Note that inequality (4.3) is better than (6.10), since the coefficient of $-n\log(n)$ in the exponent is $1$ instead of $\frac{|\lambda_{\min}|}{c}\leq 1$, avoiding the critical case when $\frac{|\lambda_{\min}|}{c}$ is very small.
Figure 3. Consider the polynomial Gaussian operator $\widehat{K}$ of the form (5.16) with parameters $A=\frac{3}{2}$, $C=1$, $B=D=E=0$, $A_{P}=-1$, $C_{P}=40$, $F_{P}=1$, and $B_{P}=D_{P}=E_{P}=0$. We plotted $q_{n}^{(i)}=q_{n,c}$ as a function of $n$ calculated from (6.2) and (4.2). For $i=1$ we obtain $c=\sum_{j}|\lambda_{j}|=1.16445$ from diagonalization, for the details see Subsection 6.2. For $i=2$ we take $c=0$, and for $i=3$ the value $c=1.4941$ was calculated from the Hölder’s upper bound and optimization given in Subsection 5.3. The horizontal dashed line indicates the common limit $-0.082228$.
Theorem 6.4.
Let $\lambda_{\min}<0$ and $\alpha=\frac{c}{|\lambda_{\min}|}$. Then for all $n$ we have
(6.9) $|q_{n,c}-\lambda_{\min}|\leq\frac{\delta_{n}}{1-\delta_{n}}|\lambda_{\min}|,$
where $10<\delta_{n}<$ 1 is the unique solution of the equation
$x^{\alpha}=2\sum_{k=n+1}^{\infty}\frac{(\alpha(1-x))^{k}}{k!}.$
For $n\geq\max\{3,2(\alpha-1)\}$ we have the estimate
(6.10) $\delta_{n}\leq\left(\frac{e\alpha}{n+1}\right)^{\frac{n+1}{\alpha}}=\exp\left[-\frac{|\lambda_{\min}|}{c}n\log(n)+\mathcal{O}(n)\right].$
Proof.
First we show (6.9). Let
$\theta=\min\{x>0:g(x)=0\}=\frac{1}{|\lambda_{\min}|}\quad\text{and}\quad\theta_{n}=\min\{x>0:h_{n}(x)=0\}.$
Since $\theta_{n}\leq\theta$, we can define $10\leq s_{n}<$ 1 such that $\theta_{n}=(1-s_{n})\theta$. Then we have
$|q_{n}-\lambda_{\min}|=\left|-\frac{1}{\theta_{n}}+\frac{1}{\theta}\right|=\frac{|\theta_{n}-\theta|}{\theta_{n}\theta}=\frac{s_{n}}{1-s_{n}}|\lambda_{\min}|,$
so as the function $x\mapsto\frac{x}{1-x}$ is monotone increasing on $[0,1)$, it is enough to show that $s_{n}\leq\delta_{n}$. For $0\leq x\leq 1$ let
$\varphi_{n}(x)=x^{\alpha}-2\sum_{k=n+1}^{\infty}\frac{(\alpha(1-x))^{k}}{k!}.$
Since $\varphi_{n}$ is strictly monotone increasing and $\varphi_{n}(\delta_{n})=0$, it is enough to prove that
(6.11) $\varphi_{n}(s_{n})\leq 0.$
First we prove that
(6.12) $g(\theta_{n})\geq s_{n}^{\alpha}.$
Let $\lambda=|\lambda_{\min}|$. It is easy to see that the function $x\mapsto(1-x)^{\frac{1}{x}}$ is monotone decreasing on $(0,1]$, so for all $0\leq x\leq\theta=\frac{1}{\lambda}$ we have
$g(x)=\prod_{i=0}^{\infty}(1+\lambda_{i}x)\geq\prod_{i:\lambda_{i}<0}(1+\lambda_{i}x)\geq\prod_{i:\lambda_{i}<0}(1-\lambda x)^{\frac{|\lambda_{i}|}{\lambda}}\geq(1-\lambda x)^{\alpha},$
which implies (6.12). On the other hand, by the definitions of $g,h_{n}$ and (3.2) for all $x\geq 0$ we obtain
$g(x)-h_{n}(x)=\sum_{k=n+1}^{\infty}\left(e_{k}+\frac{c^{k}}{k!}\right)x^{k}\leq 2\sum_{k=n+1}^{\infty}\frac{(cx)^{k}}{k!},$
therefore
(6.13) $g(\theta_{n})\leq 2\sum_{k=n+1}^{\infty}\frac{(\alpha(1-s_{n}))^{k}}{k!}.$
Then (6.12) and (6.13) imply (6.11), so the proof of (6.9) is complete.
Finally, we prove (6.10). By a well-known estimate for factorials we have
(6.14) $(n+1)!\geq\sqrt{2n\pi}\left(\frac{n+1}{e}\right)^{n+1}>4\left(\frac{n+1}{e}\right)^{n+1}$
and $n\geq\max\{3,2(\alpha-1)\}$ implies
(6.15) $\alpha(1-\delta_{n})<\alpha\leq\frac{n+2}{2}.$
Then (6.15) with Fact 4.2 and (6.14) yield
$\delta_{n}^{\alpha}=2\sum_{k=n+1}^{\infty}\frac{(\alpha(1-\delta_{n}))^{k}}{k!}\leq 4\frac{(\alpha(1-\delta_{n}))^{n+1}}{(n+1)!}\leq\left(\frac{e\alpha(1-\delta_{n})}{n+1}\right)^{n+1}.$
Substituting $N=\frac{n+1}{\alpha}$ implies
$N\delta_{n}^{\frac{1}{N}}\leq e(1-\delta_{n})\leq e.$
Thus
$\delta_{n}\leq\left(\frac{e}{N}\right)^{N}=\left(\frac{e\alpha}{n+1}\right)^{\frac{n+1}{\alpha}}=\exp\left[-\frac{|\lambda_{\min}|}{c}n\log(n)+\mathcal{O}(n)\right],$
hence (6.10) holds. This completes the proof of the theorem. ∎
6.2. Numerical calculation of the spectrum in an important special case
Matrix representation $K_{m,n}$ of an arbitrary quadratic polynomial Gaussian kernel in case of $d=1$ can be calculated rigorously, see [Newton, Equations (22),(23)]. Since the matrix elements $K_{m,n}$ decrease exponentially as $m,n\to\infty$, we can truncate the matrix at a finite size, and the original spectrum can be numerically approximated with the eigenvalues of the resulting finite dimensional operator. In the more general case, numerical diagonalization is not an easy task even for a polynomial Gaussian kernel. Our results derived from the numerical diagonalization are shown in Figure 1, and we also used it for Figure 3. The efficiency and accuracy of this method is confirmed by comparing it with the results of [Newton, balka2024positivity].
Acknowledgments
We are indebted to József Zsolt Bernád, András Bodor,
András Frigyik, Mátyás Koniorczyk, Miklós Pintér, and Géza Tóth for some helpful conversations.
Funding
The first author was supported by the National Research, Development and Innovation Office – NKFIH, grants no. 124749 and 146922, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
The second author thanks the “Frontline” Research Excellence Programme of the NKFIH (Grant no. KKP133827) and Project no. TKP2021-NVA-04, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.
The third author was supported by the Hungarian National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary grants no. 2022-2.1.1-NL-2022-00004 and 134437.
[…]